You inherited $12,000 from your Great-Uncle Michael. If you invest it and average an 8% annual rate of return (compounded annually), how much will you have in 20 years?

Show all steps.
Answer #1

Here, P = 12,000 , i = 8%, n = 20 year

Future Value = P(1 + i)n

                     = 12,000(1 + 0.08)20

                     = 12,000 * 4.661

                     = $55,932

Thus, you will have $55,932 in 20 years.